53. Comprehensive Problem (Tax Return Problem). Adam Armstrong, age 60 and single, earned $71,000 during 2018. He contributed $9,000 to the United church, paid $16,000 in federal income taxes, $3,400 in state income tax, and $4,600 in mortgage interest charges. Adam also had the following transactions: a. Sold for $60,000 stock that he had received as a gift which had an adjusted basis of $20,000 to the donor and a fair market value of $50,000 on the date of the gift and the taxable gift was $35,000 The donor had held the stock for two years and paid a gift tax of $14,000. b. Exchanged land held for five years as an investment for an apartment building and $25,000 cash. On the date of exchange, the land had an adjusted basis of $135,000. The apartment building had a fair market value of $335,000. An office building Adam owned was condemned by the state to make room for a state building. The building had an adjusted basis of $70,000 and a fair c. market value of $172,000. Adam received another building from the state and it had a fair market value of $150,000. Adam sold his personal residence for d. $145,000. He incurred $2,500 in selling expenses. The residence had an adjusted basis of $56,000. He moved into an apartment. Calculate Adams lowest taxable income. Treat all income as ordinary income.
Answer #1

Calculation of Taxable Income of Adam Armstrong:

a) If Shares are received as gift & Fair Market value exceeds Adjusted Basis then part of Gift tax paid by owner shall be added to the adjusted basis to get New basis

Here FMV = $50000 exceeds adjusted basis = $20000 then

Gift tax paid*(FMV-Adjusted Basis)/FMV = $14000*(50000-20000)/50000 = $8400

New Basis = $20000+8400 = $28400

Sale of Shares = $60000

Taxable value of Gift = 60000-28400 = $31600

b) Taxable value of Appartment on Exchange = Fair value of Appartment - Land Adjusted Basis

= $335000+25000-135000 = $225000

c) Not Taxable

d) Taxable value = sale value - selling expenses - adjusted basis = $145000-2500-56000 = $86500

Total Taxable Income = a+b+c+d+Earned Income- Contribution to Church -Mortgage Interest

= $31600+225000+86500+71000-9000-4600 = $400500